# Andreas Petrou-Zeniou

### v3

@@ Abstract
While policymakers and researchers are often concerned with conducting inference based on a data-dependent selection, a strictly larger class of inference problems arises when considering multiple data-dependent selections, such as when selecting on statistical significance or quantiles. Given this, we study the problem of conducting inference on populations selected according to their ranks, which we dub the inference on multiple winners problem. In this setting, we encounter both selective and simultaneous inference problems, making existing approaches either not applicable or too conservative. Instead, we propose a novel, two-step approach to the inference on multiple winners problem, with the first step modeling a key nuisance parameter driving selection, and the second step using this model to derive critical values on the errors of the winners. In simulations, our two-step approach reduces over-coverage error by up to 96% relative to existing approaches. In a stylized example on job training, we demonstrate that existing approaches partially apply, and that our novel two-step approach is broadly applicable and yields informative confidence sets. In a second application, we apply our two-step approach to revisit the winner’s curse in the Creating Moves to Opportunity (CMTO) program. We find that, after correcting for the inference on multiple winners problem, we fail to reject the possibility of null effects in the majority of census tracts selected by the CMTO program.
KEYWORDS: Selective Inference, Winner’s Curse, Bonferroni Inequality, Uniform Validity
JEL classification codes: C12, C13
@@ 1 Introduction
Policymakers and researchers frequently seek to conduct inferences based on a data-dependent selection. A recent paper by Andrews et al. (2023) discusses three confidence sets for a version of this problem, which they dub the inference on winners problem. However, a strictly larger class of problems arises when conducting inference on multiple such selections, such as when selecting on quantiles or cutoffs, which we dub the inference on multiple winners problem. In this setting, standard approaches either do not apply, or provide overly-conservative confidence sets. Given this, we propose a novel two-step approach to inference based on and related to the approaches from Romano et al. (2014), Canay and Shaikh (2017), and Zrnic and Fithian (2024b). In the first step, we model the selection of the multiple winners using a lower confidence bound on first-moment differences. In the second step, we use this lower confidence bound to model the error terms on the winner indices, using a Bonferroni-type correction to adjust for uncertainty in the lower confidence bound from the first step. Our goal in the inference on multiple winners problem is to conduct inference on the mean of some random vector $Y$ at the indices whose ranks according to $X$ lie in some pre-specified set $R$.
In a first application, we revisit the JOBSTART demonstration and subsequent replication studies, as in Cave et al. (1993) and Miller et al. (2005). While Cave et al. (1993) study a range of job retraining interventions for low-skilled workers at 13 separate sites, Miller et al. (2005) note that the one intervention with statistically significant effects on incomes from this original study failed to produce effects of the same magnitude in subsequent studies. We consider a thought experiment, where, instead of selecting the site with the largest treatment effect for replication, Miller et al. (2005) select all sites with statistically significant, positive treatment effects. As in Andrews et al. (2023), we study the possibility of a winner’s curse driving this replication failure. In this setting, we suggest that, because the econometrician does not know ex-ante whether one or multiple interventions will yield statistically significant treatment effects, conditional and hybrid inference procedures do not fully apply to the JOBSTART demonstration. Revised conditional and hybrid inference procedures can apply when one intervention happens to be selected, though are generally more conservative than the approaches described in Andrews et al. (2023). Moreover, depending on the significance cutoff, we find that multiple interventions can be selected with probability ranging from $17.6\%$ to $53.4\%$. This motivates our application of our two-step approach to inference in this setting. We find that our two-step approach delivers confidence sets that are 8% narrower than projection confidence sets. Our findings match those of Andrews et al. (2023) who suggest that a winner’s curse does not explain the replication failure between Cave et al. (1993) and Miller et al. (2005).
The inference on multiple winners problem is particularly relevant to the Creating Moves to Opportunity (CMTO) program from Bergman et al. (2024). Bergman et al. (2024) use the CMTO program to study the barriers to movement between low-opportunity and high-opportunity neighborhoods. In particular, Bergman et al. (2024) advertise to housing-voucher recipients selected into treatment the top third of census tracts in the Seattle commuting zone, according to the tract-level opportunity atlas developed in Chetty et al. (2018). Naturally, in selecting the top third of overall tracts (and top fifth of urban tracts) by economic mobility, an inference on multiple winners problem arises. Our novel two-step approach allows us to revisit the discussion of the CMTO program from Andrews et al. (2023), who study the aggregate effects of moving into selected, high-opportunity zones via the inference on winners setting. Our generalized setting allows us to derive simultaneous confidence sets for neighborhood effects at the tract level. Tract-level effects have previously been studied, notably in Mogstad et al. (2023). While Andrews et al. (2023) find evidence for a positive, aggregate neighborhood effects among selected tracts, Mogstad et al. (2023) show that one cannot reject the possibility that even the worst-ranked track be classified as high-opportunity in the population. However, the methods of Mogstad et al. (2023) impose simultaneous coverage on all tracts, as opposed to a subset of selected tracts. This motivates our analysis of the CMTO application. Namely, we seek to study tract-level effects at tracts labeled as high-opportunity alone. While we are able to reject the possibility of null effects at some of the tracts labeled as high-opportunity by Bergman et al. (2024), we fail to do so in 92% of high-opportunity, urban Seattle tracts. Moreover, we cannot reject the possibility of a null effect on economic mobility associated with moving from the lowest opportunity tract to an any arbitrary high-opportunity urban tract. Thus, even when carefully restricting inference to tracts of interest, we can say little about tract-level effects.
Andrews et al. (2023) provide three approaches to the inference on winners problem: conditional, projection, and hybrid. The former two approaches are described in depth in Kuchibhotla et al. (2022), who provide a review of the post-selection inference literature. Conditional approaches involve conducting inference conditional on the selection event. Lee et al. (2016) provide an exact characterization of the conditional distribution of a multivariate Gaussian conditional on polyhedral selection events, allowing for exact selective inference as in Andrews et al. (2023). Most crucially, the conditional inference procedure outlined in Andrews et al. (2023) relies on finding a univariate conditional distribution for the observed, winner’s outcome, with which they derive a confidence set via test-inversion. However, when conducting inference on multiple winners, we must compute p-values on a multivariate normal truncated to an arbitrary polyhedron, which is computationally challenging. The hybrid approach suffers from the exact same issues. The projection approach to inference, building on the work on simultaneous inference from Bachoc et al. (2017) and Berk et al. (2013), is overly conservative, since it is by design robust to arbitrary selection maps. The projection approach therefore fails to apply any ex-ante knowledge of the selection rule. Our two-step approach seeks to rectify this exact issue. We demonstrate both analytically and via simulation that the two-step approach to inference on multiple winners outperforms the projection approach. Relative to projection, the two-step approach reduces over-coverage error by up to 96.6%, and reduces confidence set volume by up to 34.5%, matching our theoretical results on the asymptotic performance of our two-step confidence sets.
@@
The paper is organized as follows: In section 2, we formally introduce the inference on multiple winners problem and present three empirically-relevant, specialized settings. In section 3, we revisit the JOBSTART demonstration and compare different approaches to inference. In section 4, we provide a more detailed discussion of the approaches to the inference on winners problem, and demonstrate that they may fail in the inference on multiple winners problem. In section 5, we construct the two-step approach to inference on multiple winners, and provide analytical results comparing its performance to the projection approach to inference. In section 6, we apply our two-step approach to inference to the CMTO program, evaluating neighborhood effects in selected census tracts. Finally, in section LABEL:Simulation_Study, we present the results from a simulation study comparing the performance of the two-step and projection approaches to inference in a range of synthetic, simulation designs. In the appendix, we provide proofs, supplemental results, and a two-step approach to an important generalization of the inference on multiple winners problem. We also compare our two-step approach to existing methods from the selective inference literature, showing favorable performance.
@@ 2 Setup and Notation
In this section, we formalize the inference on multiple winners problem. We retain the notation from Andrews et al. (2023) wherever possible. We consider a finite set of indices or populations $J:=\{1,\ldots,p\}$, the $p$-dimensional random vector $Y$, as well as a correlated random vector $X$, which are drawn jointly from a multivariate normal as below:
@@
We observe $X$ and $Y$, while $\mu$ is unknown. The variance-covariance matrix, denoted by $\Sigma$, on the other hand, is taken to be known. Our goal is to devise a valid confidence set for the values of $\mu_{Y}$ at indices selected according to the realized values of $X$. In particular, we take $R\subseteq J$ to be some fixed set of ranks and define the number of selections to be $k:=|R|$. We let the indices $j_{1},\ldots,j_{p}$ order $X$ such that $X_{j_{1}}\geq X_{j_{2}}\geq\ldots\geq X_{j_{p}}$. For example, $j_{1}$ denotes the index corresponding with the largest realization in $X$, $j_{2}$ the second largest, and so on. We define the set of selected indices $\hat{J}_{R}$ as follows:
$\hat{J}_{R}:=\{j_{l};l\in R\}$ (2)
If we take $R$ to simply be $J$, then $\hat{J}_{R}$ would be the full index set $J$. Similarly, taking $R:=\{1\}$, our selected indices would simply be the singleton $\{j_{1}\}$.
Our goal is to conduct simultaneous inference on the means $\mu_{Y,\jmath}$ for all $\jmath$ in $\hat{J}_{R}$, and in particular, construct a rectangular confidence set $CS$ in $\mathbb{R}^{k}$ indexed by $\hat{J}_{R}$ such that:
$P_{\mu,\Sigma}\left(\mu_{Y,\jmath}\in CS_{\jmath}\quad\text{for all}\quad\jmath\in\hat{J}_{R}\right)\geq 1-\alpha$ (3)
for all $\mu$ and $\Sigma$. When $R=\{1\}$, this problem is equivalent to the inference on single winners problem.
@@
Andrews et al. (2023) consider the case of inference on the winner, where $R=\{1\}$. Relatedly, Andrews et al. (2022) consider the case where selection occurs with respect to an arbitrary rank, such that $R$ is an arbitrary singleton. However, they do not consider the case where multiple rank-based selections are made. Finally, our approach easily generalizes to the case of multiple $Y$, as we discuss below.
@@ 2.1 Review of Applications
We note that, although the inference on multiple winners setting described above represents a narrow set of inference problems, it can be readily modified and generalized to incorporate a broad array of empirically-relevant settings, just as the inference on winners problem from Andrews et al. (2023). We demonstrate that many of the same applications considered in Andrews et al. (2023) remain relevant in the inference on multiple winners setting, and, additionally, a range of novel applications are also immediately tractable via inference on multiple winners approaches.
@@
Post-Selection Inference on Quantiles:  Suppose we want to conduct inference on the means of $Y$ when selecting the top $\gamma$-quantile of observations in some random vector $X$. We can take $\tau:=\lfloor\gamma p\rfloor$ and take $R=\{1,\ldots,\tau\}$. In other words, we seek to construct confidence sets for the elements in $\mu_{Y}$ corresponding to the $\tau$-best elements in $X$. This setting naturally arises in our neighborhood effects application in section 6, where we study tract-level outcomes for high opportunity tracts in the setting of Bergman et al. (2024). This specialized setting also arises in Haushofer and Shapiro (2016), who conduct a randomized controlled trial in the top 40% of villages in Rarieda, Kenya, by the proportion of thatched roofs. We may think of these proportions as drawn from a superpopulation, and seek to study superpopulation parameters at selected villages.
@@
Inference After Cutoff-Based Selections:  We may want to conduct inference on the values $\mu_{Y,j}$ for $j$ such that $X_{j}\geq c$ for some non-negative real number $c$, in the so-called file drawer problem. We can append a constant vector to our $X$ take $X_{c}=\begin{pmatrix}X^{\prime}&c\mathbb{1}_{p}^{\prime}\end{pmatrix}^{\prime}$, and $Y_{c}=\begin{pmatrix}Y^{\prime}&c\mathbb{1}_{p}^{\prime}\end{pmatrix}^{\prime}$. We can take $R:=\{1,\ldots,p\}$, and consider inference using $X_{c}$ and $Y_{c}$. The indices of the top $p$ elements in $X_{c}$ correspond exactly to those indices $j$ in $X$ such that $X_{j}\geq c$, as well as a residual set of indices corresponding (non-uniquely) to elements in the appended constant vector in $X_{c}$. [^1] The non-uniqueness in this setting is inconsequential, since those indices selected non-uniquely share a common, degenerate distribution in $Y_{c}$. In particular, the joint distribution of the $(Y_{\jmath})_{\jmath\in\hat{J}_{R}}$ is invariant to our choice of $\hat{J}_{R}$. For a concrete, empirical example, policymakers may observe the marginal values of public funds (MVPFs) of Hendren and Sprung-Keyser (2020) for a menu of policies. An MVPF exceeding one corresponds to a policy whose benefits, in dollar terms, exceeds its costs. Consequently, policymakers may choose to proceed only with policies whose MVPFs exceed one, generating an inference on multiple winners problem. While Andrews et al. (2023) consider cutoff-based selections, their framework entails performing inference on a univariate normal truncated according to a given cutoff, rather than the general case of cutoff-based selection described above. Cutoff-based selection rules have also attracted interest in the econometrics literature, namely in Gu and Koenker (2023).
@@
Inference on Statistical Significance: We notice, just as in Andrews et al. (2023), that we can normalize the $X_{j}$ by standard errors $\sqrt{\Sigma_{X,jj}}$. Since this is simply a linear transformation of the $\begin{pmatrix}X^{\prime},&Y^{\prime}\end{pmatrix}^{\prime}$, the standard inference on multiple winners setting still applies. Moreover, we can readily apply the two generalized inference on multiple winners settings as above.

### v4

@@ Abstract
This paper considers the problem of inference on multiple winners. In our setting, a winner is defined abstractly as any population whose rank according to some random quantity, such as an estimated treatment effect, a measure of value-added, or benefit (net of cost), falls in a pre-specified range of values. As such, this framework generalizes the inference on a single winner setting previously considered in Andrews et al. (2023), in which a winner is understood to be the single population whose rank according to some random quantity is highest. We show that this richer setting accommodates a broad variety of empirically-relevant applications. We develop a two-step method for inference in the spirit of Romano et al. (2014), which we compare to existing methods or their natural generalizations to this setting. We first show the finite-sample validity of this method in a normal location model and then develop asymptotic counterparts to these results by proving uniform validity over a large class of distributions satisfying a weak uniform integrability condition. Importantly, our results permit degeneracy in the covariance matrix of the limiting distribution, which arises naturally in many applications. In an application to the literature on economic mobility, we find that it is difficult to distinguish between high and low mobility census tracts when correcting for selection. Finally, we demonstrate the practical relevance of our theoretical results through an extensive set of simulations.
@@ 1 Introduction
In this paper, we consider the problem of conducting inference on multiple selections, or multiple winners. Here, we define a winner to be a population whose rank according to some random value, such as a measure of value-added or an estimated treatment effect, lies in a set specified by the analyst. We seek to provide joint confidence sets for parameters corresponding to these populations selected according to their ranks. Our framework generalizes the setting of inference on a single winner considered in Andrews et al. (2022) and Andrews et al. (2023), in that we target joint coverage of parameters among multiple selected populations. This generalization allows us to address several important applications, such as inference after cutoff-based selections, inference on quantiles, or inference on statistical significance. These settings arise frequently in applied work. For example, cutoff-based selections arise frequently in the decision-theoretic literature on optimal subset selection, as in Gu and Koenker (2023). Similarly, inference on quantiles, or equivalently inference on the $\tau$-best winners, arises in the applied literature on economic mobility, as in Bergman et al. (2024). Inference on statistical significance is particularly relevant in the literature on publication bias, as in Andrews and Kasy (2019). Our multiple winners framework identifies a common structure in these empirically relevant problems.
Motivated by this setting, we propose a novel, two-step approach to inference. In particular, we identify a key nuisance parameter that characterizes rank-based selection. In the first step of our procedure, we construct confidence bounds for this nuisance parameter. In the second step, we use these confidence bounds to bound the errors on selected units and construct critical values. We apply a Bonferroni-type correction to account for the possibility that our first step confidence region does not cover the key nuisance parameter driving selection. In this way, our approach is most similar to that of Romano et al. (2014), who study inference in moment inequality models. We demonstrate the finite-sample validity of our methods in a normal location model. We then provide results on feasible inference when the data generating process lies in a nonparametric class of distributions. We emphasize that attaining uniform validity is nontrivial in our setting, since the population selections may be indistinguishable from units not selected in a local sense. We demonstrate uniform asymptotic validity under a weak uniform integrability condition on the class of distributions generating the data. Moreover, we show that for each distribution in a large class, our two-step approach to inference asymptotically dominates projection and simultaneous inferences in the spirit of Bachoc et al. (2017), Kuchibhotla et al. (2022), and Berk et al. (2013).
Our approach lies in a broader literature on selective inference. After constructing our two-step approach to inference, we show that some well-known, existing methods for selective inference and their natural generalizations may be ill-suited for inference after selection on ranks. Projection-based methods are robust to arbitrary selection rules, and are therefore underpowered. Tools due to Lee et al. (2016) and Andrews et al. (2023) provide a polyhedral characterization of the rank-based selections considered in this paper. However, this characterization does not provide a computationally feasible inference procedure when joint coverage is desired. Consequently, it is unclear how to generalize the conditional and hybrid approaches of Andrews et al. (2023) to settings where multiple selections are made, further motivating our approach.
Within the selective inference literature, our approach is most similar in spirit to that of Zrnic and Fithian (2024b), who also apply a Bonferroni-type correction to a general selective inference setting. Similar Bonferroni-type corrections appear in Silvapulle (1996) and McCloskey (2017). As explained further in remark 3.6, the approach in Zrnic and Fithian (2024b) differs from ours, in that they propose projection inference localized to a set of likely selections, whereas our method constructs worst-case critical values over a confidence set for a key nuisance parameter driving selection. We also show analytically that our critical values are, with high probability, smaller than those of Zrnic and Fithian (2024b) for a broad set of data generating processes where both methods perform relatively poorly. The empirical relevance of this result is evident in our extensive simulations comparing the methods. In those simulation results, we also include comparisons with the alternative methodology developed in Zrnic and Fithian (2024a).
We consider two applications. In our first application, we use as an illustrative example a replication failure from two studies on job retraining due to Cave et al. (1993) and Miller et al. (2005). In our second application, we revisit the literature on economic mobility, and the problem of selecting high-opportunity census tracts in the spirit of Bergman et al. (2024). We consider the replication failure between Cave et al. (1993), who estimate treatment effects for an array of job retraining programs, and Miller et al. (2005), who replicate the study of Cave et al. (1993) for a selected subset of interventions. We build on the previous analysis of Andrews et al. (2023) to determine whether correcting for selection bias can explain the replication failure between the two studies. We find that, even when considering alternative selection rules based on statistical significance, a winners’ curse fails to explain this replication failure. In our second application, we consider the problem of ex-post inference on tract-level measures of economic mobility, as defined in Chetty et al. (2025), for a subset of high-mobility tracts considered in the randomized trial of Bergman et al. (2024). Our analysis revisits the findings of Mogstad et al. (2023), who suggest that the selection of high-opportunity tracts may reflect noise as opposed to signal. We find similar results to Mogstad et al. (2023), in that we generally fail to reject the null that pairwise differences between arbitrary high and low opportunity tracts are zero. We obtain more positive results, however, in comparing commuting zones.
Finally, we conduct an extensive simulation study comparing our methods to existing methods in the literature, namely projection inference as defined in Andrews et al. (2023), the locally simultaneous approach of Zrnic and Fithian (2024b), and a recent approach due to Zrnic and Fithian (2024a). Our methods are able to outperform these existing methods across a broad range of simulation designs, both in terms of reducing over-coverage and in reducing confidence set length. In particular, our methods reduce over-coverage error by up to 96%, and reduce confidence set length by up to 27% relative to projection inference. We also reduce over-coverage error by up to 71% and length by up to 11% relative to locally simultaneous inference.
@@
The paper is organized as follows: In section 2, we formally introduce the inference on multiple winners problem and present four empirically-relevant settings to which it applies. In section 3, we develop the two-step approach to inference on multiple winners. In section 4, we provide a more detailed discussion of existing approaches to the inference on winners problem, outline some of their shortcomings in the inference on multiple winners problem, and provide some analytical results comparing our two-step approach to inference to these existing methods. In section 5, we revisit the JOBSTART demonstration and compare different approaches to inference. In section 6, we apply our two-step approach to inference to the CMTO program, evaluating neighborhood effects in selected census tracts. Finally, in section 7, we present the results from a simulation study comparing the performance of the two-step and projection approaches to inference in a range of synthetic, simulation designs. In the appendix, we describe further approaches in the literature, provide proofs and supplemental results, and provide a full simulation study.
@@ 2 Setup and Notation
In this section, we formalize the inference on multiple winners problem. We consider a finite set of indices or populations $J:=\{1,\ldots,p\}$, the $p$-dimensional random vector $Y$, as well as a correlated random vector $X$, which are drawn jointly from a multivariate normal as below:
@@
We observe $X$ and $Y$, while $\mu$ is unknown. The variance-covariance matrix, denoted by $\Sigma$, on the other hand, is taken to be known. Our goal is to devise a valid confidence set for the values of $\mu_{Y}$ at indices selected according to the realized values of $X$. In particular, we take $R\subseteq J$ to be some fixed set of ranks of interest and define the number of selections to be $k:=|R|$. We let the indices $j_{1}(X),\ldots,j_{p}(X)$ order $X$ such that $X_{j_{1}(X)}\geq X_{j_{2}(X)}\geq\ldots\geq X_{j_{p}(X)}$. For example, $j_{1}(X)$ denotes the index corresponding to the largest realization in $X$, $j_{2}(X)$ the second largest, and so on. We define the set of selected indices $\hat{J}_{R}(X)$ as follows:
$\hat{J}_{R}(X):=\{j_{l}(X);l\in R\}$ (2)
If we take $R$ to simply be $J$, then $\hat{J}_{R}(X)$ would be the full index set $J$. Similarly, taking $R:=\{1\}$, our selected indices would simply be the singleton $\{j_{1}(X)\}$. Unless stated otherwise, we will denote $\hat{J}_{R}(X)$ by $\hat{J}_{R}$ in the following analyses. While $j_{l}(X)$ need not be unique for $l$ in $R$, the methods presented in this paper are robust to arbitrary tie-breaking rules.
Our goal is to conduct simultaneous inference on the means $\mu_{Y,\hat{\jmath}}$ for all $\hat{\jmath}$ in $\hat{J}_{R}$, and in particular, construct a rectangular confidence set $CS$ in $\mathbb{R}^{k}$ indexed by $\hat{J}_{R}$ such that:
$P_{\mu,\Sigma}\left(\mu_{Y,\hat{\jmath}}\in CS_{\hat{\jmath}}\quad\text{for all}\quad\hat{\jmath}\in\hat{J}_{R}\right)\geq 1-\alpha$ (3)
for all $\mu$ and $\Sigma$. When $R=\{1\}$, this problem is equivalent to the inference on (single) winners problem considered in Andrews et al. (2023). When convenient, we may write $(\mu_{Y,j})_{j\in J_{c}}\in CS$ to denote $\mu_{Y,j}\in CS_{j}\text{ for all }j\in J_{c}$, where $J_{c}$ is an arbitrary subset of $J$.
@@
In addition to the case of $R=\{1\}$ considered in Andrews et al. (2023), Andrews et al. (2022) consider the case where selection occurs with respect to an arbitrary rank, such that $R$ is an arbitrary singleton. However, they do not consider the case where multiple rank-based selections are made.
@@ 2.1 Review of Applications
Before proceeding, we describe a broad array of empirically-relevant settings in which our results are of interest. In addition to the same applications considered in Andrews et al. (2023), the inference on multiple winners setting accommodates a range of novel applications.
@@
Post-Selection Inference on Quantiles:  Suppose we want to conduct inference on the components of the mean of $Y$ corresponding to the components of $X$ in the top $\gamma$-quantile of all components of $X$. We can take $\tau:=\lfloor\gamma p\rfloor$ and take $R=\{1,\ldots,\tau\}$. In other words, we seek to construct confidence sets for the elements in $\mu_{Y}$ corresponding to the $\tau$-best elements in $X$. This setting naturally arises in our neighborhood effects application in section 6, where we study tract-level outcomes for high opportunity tracts in the setting of Bergman et al. (2024). This specialized setting also arises in Haushofer and Shapiro (2016), who conduct a randomized controlled trial in the top 40% of villages in Rarieda, Kenya, selected according to the proportion of houses with thatched roofs.
@@
Inference After Cutoff-Based Selections:  We may want to conduct inference on the values $\mu_{Y,j}$ for $j$ such that $X_{j}\geq c$ for some non-negative real number $c$, in the so-called file drawer problem. In order to accommodate this setting in our framework, define $X_{c}:=\begin{pmatrix}X^{\prime}&c\mathbb{1}_{p}^{\prime}\end{pmatrix}^{\prime}$, and $Y_{c}:=\begin{pmatrix}Y^{\prime}&c\mathbb{1}_{p}^{\prime}\end{pmatrix}^{\prime}$. We can take $R:=\{1,\ldots,p\}$, and consider inference using $X_{c}$ and $Y_{c}$. The indices of the top $p$ elements in $X_{c}$ correspond exactly to those indices $j$ in $X$ such that $X_{j}\geq c$, as well as a residual set of indices corresponding (non-uniquely) to elements in the appended constant vector in $X_{c}$. [^1] The non-uniqueness in this setting is inconsequential, since those indices selected non-uniquely share a common, degenerate distribution in $Y_{c}$. In particular, the joint distribution of the $(Y_{\hat{\jmath}})_{\hat{\jmath}\in\hat{J}_{R}}$ is invariant to our choice of $\hat{J}_{R}$. For a concrete, empirical example, policymakers may observe the marginal value of public funds (MVPFs) of Hendren and Sprung-Keyser (2020) for a menu of policies. An MVPF exceeding one corresponds to a policy whose benefits, in dollar terms, exceeds its costs. Consequently, policymakers may choose to proceed only with policies whose MVPFs exceed one, generating an inference on multiple winners problem. Our analysis of cutoff-based selections differs from that of Andrews et al. (2023) in that we can accommodate multiple selections. In contrast, the setting of Andrews et al. (2023) can only accommodate the above when $p=1$. Cutoff-based selection rules have also attracted interest in the literature on optimal selection, namely in Gu and Koenker (2023).
@@
Inference on Statistical Significance: We observe, just as in Andrews et al. (2023), that we can normalize the $X_{j}$ by standard errors $\sqrt{\Sigma_{X,jj}}$. Since this is simply a linear transformation of the $\begin{pmatrix}X^{\prime},&Y^{\prime}\end{pmatrix}^{\prime}$, the standard inference on multiple winners setting still applies. By applying the previous two examples, we can accommodate inference on all units that are statistically significant at a given level according to $X$. Note that $X$ and $Y$ need not be the same. Such problems arise in the literature on publication bias, as in Andrews and Kasy (2019).
Inference on Multiple Outcomes: We may be concerned with the means of several outcomes of interest among selected populations. For example, in the CMTO application of Bergman et al. (2024), we may be concerned with mobility effects among different groups (say effects by race or gender). Denote by $Y_{1},\ldots,Y_{K}$ the different outcomes of interest. In order to accommodate this setting in our framework, define $Y_{r}:=\begin{pmatrix}Y_{1}^{\prime}&\ldots&Y_{K}^{\prime}\end{pmatrix}^{\prime}$ and $X_{r}:=\begin{pmatrix}X^{\prime}&\ldots&X^{\prime}\end{pmatrix}^{\prime}$, where $X$ is repeated $K$ times. Further define a new index set $R_{r}:=\{l\cdot q;l\in R,\ 1\leq q\leq K\}$. The quantities $Y_{r}$, $X_{r}$, and $R_{r}$ defined in this way characterize the inference on multiple winners problem with the desired estimands.
Remark 2.2.
In many of these examples of applications, $\Sigma$ need not be full rank. Our methods are valid in finite samples when imposing normality as in (1). Moreover, when deriving asymptotic properties of our methods, we demonstrate we can provide uniformly valid inferences over a class of data generating processes which may included distributions with degenerate covariance matrices. We discuss and provide necessary conditions for asymptotic validity in appendix B.
3 A Two-Step Approach to Inference on Multiple Winners
In this section, we construct a two-step approach to inference on multiple winners. We construct a confidence interval that is conceptually similar to the inference approach outlined Romano et al. (2014), and most recently in the selective inference literature, to Zrnic and Fithian (2024b). We emphasize, however, that our approach is meaningfully different from that of Zrnic and Fithian (2024b); see remark 3.6 for further discussion. In the first step, we construct confidence bound on moment differences. In the second step, we use these bounds to model the errors on selected units and derive critical values. As we will show in sections 4 and 7, our approach to inference performs well relative to existing methods.
3.1 Construction
Our goal, as usual, is to construct a rectangular confidence set $CS_{1-\alpha}^{TS}$ satisfying (3), where $CS_{1-\alpha}^{TS}=\times_{j\in\hat{J}_{R}}CS_{1-\alpha,j}^{TS}$. Our construction is based on the following observation, which states that:
$\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}\leq\max_{j\in J}\frac{|\xi_{Y,j}|}{\sqrt{\Sigma_{Y,jj}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+\Delta_{j^{\prime}j}\right)\in R\right)$ (4)
for all $\hat{\jmath}\in\hat{J}_{R}$, where $\Delta_{jj^{\prime}}:=\mu_{X,j}-\mu_{X,j^{\prime}}$ is an unobserved nuisance parameter. Our goal is therefore to construct a confidence region given by a lower bound $L$ and upper bound $U$ for $\Delta$, such that $P_{\mu,\Sigma}\left(L\leq\Delta\leq U\right)\geq 1-\beta$ for some choice of $\beta$ in $(0,\alpha)$, where the inequality is interpreted element-wise. We use this lower bound to construct a probabilistic upper bound for the right-hand side of (4). We finally use a Bonferroni-type correction to correct for the possibility that the event $\{L\leq\Delta\leq U\}$ does not hold.
We construct $L$ by first taking $d_{1-\beta/2}(\Sigma)$ to be the $1-\beta/2$-quantile of the following:
$\max_{j,j^{\prime}\in J,j\neq j^{\prime},\text{var}_{jj^{\prime}}\neq 0}\frac{|\xi_{X,j}-\xi_{X,j^{\prime}}|}{\sqrt{\text{var}_{jj^{\prime}}}}~,$ (5)
where $\text{var}_{jj^{\prime}}:=\text{Var}(\xi_{X,j}-\xi_{X,j^{\prime}})$. We define $L$ and $U$ as follows:
$L_{jj^{\prime}}$ $:=X_{j}-X_{j^{\prime}}-d_{1-\beta/2}(\Sigma)\sqrt{\text{var}_{jj^{\prime}}}$ (6)
$U_{jj^{\prime}}$ $:=X_{j}-X_{j^{\prime}}+d_{1-\beta/2}(\Sigma)\sqrt{\text{var}_{jj^{\prime}}}~.$ (7)
It follows that $P_{\mu,\Sigma}\left(L\leq\Delta\leq U\right)\geq 1-\beta$. We note that, for each $\hat{\jmath}\in\hat{J}$, the following inequality holds:
$\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}\leq f(L,U)$ (8)
on the event $B:=\{L\leq\Delta\leq U\}$, where, for non-random vectors $\ell$ and $u$,
$f(\ell,u):=\max_{j\in J}\frac{|\xi_{Y,j}|}{\sqrt{\Sigma_{Y,jj}}}\mathbb{1}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+u_{j^{\prime}j}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+\ell_{j^{\prime}j}\right)\right]\cap R\neq\emptyset\right)~.$
Denote by $\rho_{1-\alpha+\beta}(L,U)$ the $1-\alpha+\beta$ quantile of the $f(\ell,u)$, evaluated at the random values of $L$ and $U$, and consider the following, rectangular confidence interval:
$CS_{1-\alpha;\beta}^{TS}:=\times_{\hat{\jmath}\in\hat{J}_{R}}\left[Y_{j}-\rho_{1-\alpha+\beta}(L,U)\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}},Y_{j}+\rho_{1-\alpha+\beta}(L,U)\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}\right$]
We claim the following:
Proposition 3.1.
$CS_{1-\alpha;\beta}^{TS}$ is a valid confidence set at the $(1-\alpha)$-level, such that:
$P_{\mu,\Sigma}\left(\left(\mu_{Y,\hat{\jmath}}\right)_{\hat{\jmath}\in\hat{J}_{R}}\in CS_{1-\alpha;\beta}^{TS}\right)\geq 1-\alpha$ (9)
$P_{\mu,\Sigma}\left(\left(\mu_{Y,\hat{\jmath}}\right)_{\hat{\jmath}\in\hat{J}_{R^{\prime}}}\in CS_{1-\alpha;\beta}^{TS}\right)\geq 1-\alpha$ (10)
for all $\mu$, $\Sigma$.
We prove this result in appendix C.
Remark 3.1.
In appendix A, we construct a version of our two-step confidence set that is intersected with the projection confidence set for both $\mu_{X}$ and $\mu_{Y}$, where we formally define the projection confidence set in section 4.2. Informally, we can think of the projection confidence set $CS_{1-\alpha}^{P}$ as a special case of the two-step confidence set with $\beta=0$, and with $L=$-$\infty$ and $U=\infty$, where here $\pm\infty$ are interpreted as vectors. Notably, when $X=Y$ it follows that this intersected two-step confidence set is a subset of projection. This case corresponds to the case where units are selected according to their ranks in $Y$, or equivalently, where $\hat{J}_{R}(X)=\hat{J}_{R}(Y)$.
Remark 3.2.
In the case where $R$ is of the form $\{1,\ldots,\tau\}$, we can simplify our two-step approach to inference. In particular, we can write (4) as follows:
$\frac{|\xi_{Y,\hat{\jmath}}|}{\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}}\leq\max_{j\in J}\frac{|\xi_{Y,j}|}{\sqrt{\Sigma_{Y,jj}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+\Delta_{j^{\prime}j}\right)\geq p-\tau\right).$
Taking $\tilde{L}_{jj^{\prime}}:=X_{j}-X_{j^{\prime}}-d_{1-\beta}(\Sigma)\sqrt{\text{var}_{jj^{\prime}}}$, and defining $\tilde{B}:=\left\{\tilde{L}\leq\Delta\right\}$, we have that on the event $\tilde{B}$: